You borrow $5,000 from your parents to purchase a used car. The arrangements of the loan are such that you make payments of $250 per month toward the balance plus 1% interest on the unpaid balance from the previous month. (a) Find the first year’s monthly payments and the unpaid balance after each month. (b) Find the total amount of interest paid over the term of the loan
Answer by Theo(13342)

to completely satisfy this loan, it would take approximately 22.43 months, at which time the remaining balance due will be 0.

this calculation was performed by using a time value of money financial calculator at https://arachnoid.com/finance/

the results of the calculation are shown below:

$$$

you enter:

present value = 5000 (positive because this is money that comes in to you)
future value = 0
number of periods = 0 or blank since this is what you want to calculate.
payment = -250 (negative because this is money that goes out from you)
interest rate per time period = 1%
payment is made at the end of each month.

you then hit the NP button and the calculator tells you that the number of months required to make the remaining balance equal to 0 is 22.43.

this is a rounded number.
the actual number that is less rounded, or at least rounded to more decimal digits, is equal to 22.42574188.

i used another financial calculator to get that number.

the month by month calculations are shown in the following excel spreadsheet.

$$$

the first year's monthly payments are equal to 12 * 250 = $3000.00.

the total interest paid is equal to $606.74 which can be calculated by summing up the interest payments for each month from the excel spreadsheet which i did for you.

the other way of calculating the interest payments would be to sum up all the monthly payments and then subtracting the original loan from that.

for example, the original loan was 5000 and the sum of the total monthly payments for 22.42574188 month is equal to 22.42574188 * 250 = $5606.43547.
subtract 5000 from that and you get $606.44 rounded to 2 decimal places.

the spreadsheet shows $607.74.

the difference is most likely due to rounding.

if you go by the results of the online calculator, then the total interest would be 250 * 22.43 minus 5000 = $607.5.

i would go with that because you have a reference that you can point to it there's any question.

you can also calculate by formula but i ran out of time to derive the formula for you.

i will do it later if you require it, but not now.

my recommendation.

use the calculator that you have available to determine the number of months and then use that to determine the total interest payments.

let me know if you have any further questions regarding this.

i'll answer by tomorrow morning at the latest.